can anyone help me to find this answer??
a production editor at a textbook publishing house decided that the pages of abook should have 1-inch margins at the top and bottom and 0.5-inch margins on the side. she further stipulated that the length of a page should be 1.5 times its width and have a printed area of exactly 51 square inches.find the dimensions of a page of the book
thank you...

Answer by mananth(16946)

let the width be x
the length will be 1.5x
..
Side Margin = 0.5 inch. 1 " total
Top margin 1 " bottom 1" = 2 inches.
Printing area = (x-1)(1.5x-2)=51
1.5x^2-2x-1.5x+2=51
1.5x^2-3.5x+2=51
multiply by 2
3x^2-7x+4=102
3x^2-7x-98=0
3x^2-21+14x-98=0
3x(x-7)+14(x-7)=0
(x-7)(3x+14)
x= 7 ignore -14/3
printing dimensions = 7" & 10.5 "